The State of Chiapas, Mexico, decided to fund a program for improving reading skills in elementary school students. The first cost is $325,000 now and an update amount of $160,000 every 5 years forever. Determine the perpetual equivalent annual cost at an interest rate of 11% per year.

The perpetual equivalent annual cost is determined to be –$
Answer #1

PW = -P - [F(A/F, i, n)] / i

       = -325,000 - [160,000(A/F, 11%, 5)] / 0.11

       = -325,000 - [160,000(0.1606)] / 0.11

       = -325,000 - 233,600

       = -558,600

A = -558,600(A/P, 11%,∞)

   = -558,600 * 0.11

   = -$61,446

The perpetual equivalent annual cost is determined to be -$61,446